What is the diameter of the circle circumscribing an isosceles right triangle having an area of 126 sq.cm?


1
Expert's answer
2021-05-13T17:47:31-0400

Ans:- Given that an isosceles right triangle having an area of 126 sq.cm.


"\Rightarrow" Area of right angled isosceles triangle "=\dfrac{1}{2}\times a\times a=126\to a=15.87cm"


From Pythagorean Theorem , "c^2=a^2+b^2 \to c^2=2a^2" ("a=b" because of isosceles triangle )


"c=1.414\times a \to c=1.414\times15.87=22.443cm"


So, diameter of circle "=22.443cm"